# Solving Linear Systems of Equations with the Quantum HHL Algorithm: A Tutorial on the Physical and Mathematical Foundations for Undergraduate Students

### v1

@@ I Introduction
Quantum computing has gained prominence in recent years since the proposal of constructing a quantum computer by Feynman in a series of seminars held in 1981 Feynman (2018). More recently, with the proposal of quantum algorithms capable of outperforming classical algorithms, such as Grover’s algorithm for searches in an unstructured list Grover (1996) or Shor’s algorithm for integer factorization Shor (1994), the field has gained industrial attention with billions of dollars invested annually Castelvecchi (2025).
@@
Also of great importance in the field is the quantum algorithm for solving systems of linear equations (SLE) proposed by Harrow, Hassidim, and Lloyd in 2009 Harrow et al. (2009), widely known as HHL. While classically the execution complexity of these algorithms can be on the order of $O(N^{3})$, HHL demonstrates exponential acceleration, performing the same task with complexity $poly(\log N)$, where $N$ is the number of linear equations Harrow et al. (2009); Zaman et al. (2023). The algorithm proposes solving SLEs in such a way as to obtain an approximate or proportional value to the solution vector, using methods widely covered in textbooks, such as state preparation, quantum phase estimation, and controlled rotation Nielsen and Chuang (2010); Mermin (2007). In the field of applied quantum computing research, the algorithm has gained prominence in various applications, for example, quantum machine learning Biamonte et al. (2017); Duan et al. (2020a); Das Sarma et al. (2019), quantum finance Rebentrost and Lloyd (2024), differential equation solving Xin et al. (2020); Subaşı et al. (2019), linear regression Nielsen and Chuang (2010), and, more recently, cryptosystem attacks Chen and Gao (2022).
@@
However, the vast majority of these works focus on the application of the algorithm, assuming that the reader is already familiar with its theoretical foundations Biamonte et al. (2017); Das Sarma et al. (2019); Rebentrost and Lloyd (2024); Xin et al. (2020); Subaşı et al. (2019); Nielsen and Chuang (2010); Chen and Gao (2022). Although in recent years the field of quantum computing and information has attracted the attention of undergraduate students and the scientific community Rabelo and Costa (2018); Jesus et al. (2021); Canabarro et al. (2022); Fernandes et al. (2022); Alves and Felipe (2022); Alves et al. (2020), materials that describe in detail the characteristics of the HHL algorithm are still scarce Zaman et al. (2023); Duan et al. (2020a); Zhang et al. (2022), especially in Portuguese Galvão et al. (2024). In this context, the literature still lacks accessible resources to assist beginners in understanding and applying more complex quantum algorithms, such as undergraduate students in Physics and Computer Science.
@@
To address this gap, we present in this work a tutorial with the main physical and mathematical foundations necessary for the application of the HHL algorithm in solving systems of linear equations. In this sense, the aim of the text is to provide an overview of the algorithm, discussing its physical and mathematical foundations along with guided implementation exercises in Qiskit, ensuring that the reader understands both the theory and its practical application to an SLE. Thus, this material can be used as a tutorial for students and instructors in Physics and Computer Science, since it emerges at the interface between these fields of knowledge. We therefore hope to contribute to the Physics Education landscape by making the details of the algorithm accessible to undergraduates, aligning with other materials already available in the field Galvão et al. (2024); Jesus et al. (2021); Billig (2018); Oliveira et al. (2021); Santos (2017); Rabelo and Costa (2018); Alves et al. (2020); José et al. (2013); Perry et al. (2019); Oliveira (2020).
@@ II HHL Algorithm
In this section, we derive the mathematical foundations of the algorithm, detailing algebraically how the application of specific circuits affects the result. If necessary, a general introduction to quantum computing can be found in Appendix A. We also recommend the works Galvão et al. (2024); Jesus et al. (2021); Billig (2018); Oliveira et al. (2021); Santos (2017); Rabelo and Costa (2018); Alves et al. (2020); José et al. (2013); Perry et al. (2019); Oliveira (2020). Additionally, we suggest materials that may help the reader with specific steps, such as state preparation Aulicino et al. (2022) and quantum phase estimation Alves and Felipe (2022); Pezze and Smerzi (2014).
@@ II.1 Linear Systems of Equations
Linear Systems of Equations (SLE) are fundamental for solving problems in both science and industry. They are widely used to model physical phenomena, such as current flow in electrical circuits, force equilibrium in mechanical structures, and mesh simulation in numerical methods Lanczos (1952); Oktaç (2018); Crout (1941); Kirvan et al. (2015). Furthermore, SLEs underpin algorithms in machine learning, optimization, and image processing, and are often solved at large scale on supercomputers Cleveland Ashcraft et al. (1987); Saad (1989). Thus, developing efficient methods for solving these systems, such as the HHL algorithm in quantum computing, represents an important step toward accelerating scientific and industrial applications.
@@
Here, $A$ is an $N\times N$ matrix and $\vec{x}$ and $\vec{b}$ are $N$-dimensional vectors. In this paper, we adopt the usual notation from Quantum Mechanics, using bras $\left\langle\cdot\right|$ and kets $\left|\cdot\right\rangle$, so that Eq. (1) can be rewritten as $A\left|x\right\rangle=\left|b\right\rangle$ (see Appendix A). Notoriously, the greatest difficulty in solving such systems lies in inverting the matrix $A$ to obtain the solution vector Diblík et al. (2016):
@@
For very large values of $N$, this process requires significant computational resources, since classically the complexity of the problem can scale polynomially Zaman et al. (2023). HHL promises to reduce this complexity to logarithmic scale, allowing SLEs to be solved with fewer computational resources Harrow et al. (2009).
@@ II.2 General Overview of the Algorithm
The HHL algorithm uses in its circuit steps that are widely known in the Quantum Computing literature Nielsen and Chuang (2010). Figure 1 shows a general overview of the circuit, explicitly illustrating its routines and applied registers. In summary, in order of implementation, the steps of the algorithm can be synthesized as:
@@ II.4 Quantum Phase Estimation
Quantum Phase Estimation (QPE) was introduced in 1995 by Alexei Kitaev to estimate the eigenvalue phase of an eigenvector associated with a unitary operator Kitaev (1995). Considering this unitary operator as a matrix $\mathbf{U}$, with eigenvectors $\left|u\right\rangle$, its characteristic equation can be written as
@@ II.4.1 Uniform Superposition
The first step of QPE is similar to the initial stage present in many quantum algorithms: creating a uniformly distributed superposition of states. A famous example of this application is Grover’s algorithm, in which all qubits are initialized this way Grover (1996). In the case of QPE, this superposition is applied only to the registers $q-c:\left|0\right\rangle^{\otimes n}$. This relation can be obtained by applying the Hadamard gate ($H$) to these qubits:
@@ II.4.2 Controlled Unitary Operators
The sequence of products in (16), directly related to $l$ in Eq. (15), is given by the number of qubits that form the qcq-c registers. This number is chosen by the user and defines the precision of the phase calculated by QPE. Using more qubits increases precision, but care must be taken not to incur excessive computational cost for very small changes Papadopoulos et al. (2024). Moreover, larger circuits are more susceptible to errors due to the current structure of quantum computers. Present-day machines have an intermediate number of physical qubits (between tens and hundreds) and suffer from error and noise rates that prevent the implementation of error correction codes, which characterizes them as NISQ devices (Noisy Intermediate-Scale Quantum) Preskill (2018). Computational complexity is also a contributing factor in choosing fewer qubits: making the system too complex may waste resources and runtime. A more detailed analysis of the impact of the number of qubits on QPE precision can be found in Ref. Koch et al. (2020).
@@
Equation (17) contains a relation known as the Quantum Fourier Transform (QFT), an algorithm used to switch between spaces in quantum systems Nielsen and Chuang (2010); Coppersmith (1994). Its general expression, for two arbitrary states $\left|\alpha\right\rangle$ and $\left|\beta\right\rangle$, can be highlighted as:
@@ II.4.3 Inverse Quantum Fourier Transform
which occurs only when $\theta-y/\tilde{N}=0$, where $\tilde{N}=2^{n}$. This result implies that in any other situation the expression cancels out, since the Fourier transform is based on the sum of sinusoids and, like the Dirac delta function, has a filtering property Zaman et al. (2023).
@@ II.5 Ancilla Quantum Encoding
Ancilla qubits are widely used in quantum algorithms, as they assist in specific implementation steps such as indirect measurements García-Pérez et al. (2021); Brida et al. (2012), state control Brown et al. (2011); Zhang and Sachdev (2020), or error correction Reinhold et al. (2020); Criger et al. (2012). In the context of HHL, this qubit acts to maximize the probability of obtaining the desired result, given the eigenvalues encoded in the clock qubits during QPE Harrow et al. (2009). Without loss of generality, one can write a controlled rotation on the state as
@@
Previous works define specific functions to obtain appropriate angles when applying rotations with the gate $R_{Y}(\vec{\theta})$ Zaman et al. (2023); Zhang et al. (2022), namely
@@ III Numerical Implementation
Using these values to determine the corresponding rotations in the quantum circuit will be discussed in later subsections. However, this is not a mandatory requirement of the algorithm but rather an efficient didactic choice to understand the theoretical foundations discussed in Sec. II. In this way, one can visualize how specific rotations influence the final result. This is because the HHL algorithm does not require prior knowledge of the eigenvalues and eigenvectors of the matrix $A$; such information is accessed indirectly through quantum phase estimation, provided the matrix is sparse and well-conditioned, and an efficient oracle exists to simulate $e^{iAt}$. Thus, in cases where the matrix $A$ is unitary and can be implemented via quantum logic gates, the appropriate choice of rotations can be made efficiently, without the need for explicit diagonalization Harrow et al. (2009).
In the following subsections, we present code for implementing the HHL routines using Qiskit, IBM’s open-source framework for programming and simulating quantum circuits on real hardware and simulators; the code can be found in Galvão et al. (2025). The implementation is carried out in Python, requiring basic familiarity with the language. For an introduction to Python, we recommend dos Santos Cruz et al. (2022); Borges (2014).
@@ III.2 Quantum Phase Estimation
$\begin{split}\left|\varphi_{2}\right\rangle&=\left|1\right\rangle\frac{1}{2}(\left|00\right\rangle+\left|01\right\rangle+\left|10\right\rangle+\left|11\right\rangle)\left|0\right\rangle_{a},\\ \left|\varphi_{2}\right\rangle&=\frac{1}{\sqrt{2}}(-\left|u_{1}\right\rangle\left|00\right\rangle-\left|u_{1}\right\rangle\left|01\right\rangle-\left|u_{1}\right\rangle\left|10\right\rangle-\left|u_{1}\right\rangle\left|11\right\rangle\\ &+\left|u_{2}\right\rangle\left|00\right\rangle+\left|u_{2}\right\rangle\left|01\right\rangle+\left|u_{2}\right\rangle\left|10\right\rangle+\left|u_{2}\right\rangle\left|11\right\rangle)\,{\left|0\right\rangle_{a}}.\end{split}$ (44)
@@
$\mathbf{U}=\frac{1}{2}\begin{pmatrix}-1+i&1+i\\ 1+i&-1+i\\ \end{pmatrix}$. (48)
Obtaining higher powers of the unitary reduces to $\mathbf{U^{l}}=V\mathbf{U_{d}^{l}}V^{-1}$. For $n=2$ c-qubits, $l=2$ is sufficient to provide acceptable precision for the algorithm (QPE is a phase estimation and using more qqcc qubits refines this estimate; the measured value gets closer to the expected phase, but the measurement probability may decrease Koch et al. (2020)). Thus, $\mathbf{U^{2}}$ is written as
@@ III.5 Numerical Results
To further evaluate the algorithm’s potential on NISQ-era hardware Preskill (2018), we also executed it on a real quantum computer, ibm_kiyv, available via the IBM Quantum Experience; its specifications are shown in Tab. 1. For a more reliable metric, we ran the algorithm $10$ times on this backend and computed the mean and standard deviation. Both results are shown in Fig. 5, with the probability distributions produced by the algorithm.
@@ IV Noise Analysis in the Algorithm
In the context of the NISQ era, several factors can affect the efficiency of the algorithm, such as gate errors, relaxation noise ($T_{1}$), dephasing noise ($T_{2}$), and errors inherent to the measurement process Preskill (2018). Each of these noise sources contributes in different ways to information loss and to the deviation between ideal and experimentally obtained results.
@@
This behavior highlights the importance of precise control over two-qubit gates—which are notoriously noisier than single-qubit operations—as well as the need to minimize circuit depth to reduce decoherence effects. To mitigate the errors observed on real quantum hardware, several strategies can be applied. One of the most direct approaches is readout error mitigation, which calibrates measurement probabilities using known reference states, enabling a more accurate reinterpretation of the observed outcomes Hicks et al. (2022); Smith et al. (2021); Yang et al. (2022).
@@
In more advanced scenarios, implementing quantum error-correcting codes such as the surface code may be considered, although they still require a large number of physical qubits to protect a single logical qubit, which makes them impractical for many present-day systems Wang et al. (2011); Horsman et al. (2012); Hill et al. (2015). Finally, statistical repetition and post-processing techniques are also useful for smoothing noise effects, especially when combined with theoretical modeling Maciejewski et al. (2020); Zhukov and Pogosov (2022). Simulation with noise models, as presented in this work, also allows one to anticipate behaviors and adjust strategies before running on hardware, representing an essential step toward developing more robust quantum algorithms.

### v2

@@ I Introduction
Quantum computing has gained prominence in recent years since the proposal of constructing a quantum computer by Feynman in a series of seminars held in 1981 feynman2018simulating. More recently, with the proposal of quantum algorithms capable of outperforming classical algorithms, such as Grover’s algorithm for searches in an unstructured list grover1996fast or Shor’s algorithm for integer factorization shor1994algorithms, the field has gained industrial attention with billions of dollars invested annually Castelvecchi2025.
@@
Also of great importance in the field is the quantum algorithm for solving systems of linear equations (SLE) proposed by Harrow, Hassidim, and Lloyd in 2009 harrow2009quantum, widely known as HHL. While classically the execution complexity of these algorithms can be on the order of $O(N^{3})$, HHL demonstrates exponential acceleration, performing the same task with complexity $poly(\log N)$, where $N$ is the number of linear equations harrow2009quantum; zaman. The algorithm proposes solving SLEs in such a way as to obtain an approximate or proportional value to the solution vector, using methods widely covered in textbooks, such as state preparation, quantum phase estimation, and controlled rotation nielsen2010quantum; mermin2007quantum. In the field of applied quantum computing research, the algorithm has gained prominence in various applications, for example, quantum machine learning biamonte2017quantum; duan2020survey; sarma2019machine, quantum finance rebentrost2018quantum, differential equation solving xin2020quantum; subacsi2019quantum, linear regression nielsen2010quantum, and, more recently, cryptosystem attacks chen2022quantum.
@@
However, the vast majority of these works focus on the application of the algorithm, assuming that the reader is already familiar with its theoretical foundations biamonte2017quantum; sarma2019machine; rebentrost2018quantum; xin2020quantum; subacsi2019quantum; nielsen2010quantum; chen2022quantum. Although in recent years the field of quantum computing and information has attracted the attention of undergraduate students and the scientific community rabelo2018abordagem; jesus2021computaccao; canabarro2022quantum; fernandes2022ions; alves2022algoritmos; alves2020simulating, materials that describe in detail the characteristics of the HHL algorithm are still scarce zaman; duan2020survey; zhang2022improved, especially in Portuguese galvao2024possibilidades. In this context, the literature still lacks accessible resources to assist beginners in understanding and applying more complex quantum algorithms, such as undergraduate students in Physics and Computer Science.
@@
To address this gap, we present in this work a tutorial with the main physical and mathematical foundations necessary for the application of the HHL algorithm in solving systems of linear equations. In this sense, the aim of the text is to provide an overview of the algorithm, discussing its physical and mathematical foundations along with guided implementation exercises in Qiskit, ensuring that the reader understands both the theory and its practical application to an SLE. Thus, this material can be used as a tutorial for students and instructors in Physics and Computer Science, since it emerges at the interface between these fields of knowledge. We therefore hope to contribute to the Physics Education landscape by making the details of the algorithm accessible to undergraduates, aligning with other materials already available in the field galvao2024possibilidades; jesus2021computaccao; billig2018quantum; oliveira2021algoritmos; santos2017computador; rabelo2018abordagem; alves2020simulating; jose2013introduccao; perry2019quantum; oliveira2020fisica.
@@ II HHL Algorithm
In this section, we derive the mathematical foundations of the algorithm, detailing algebraically how the application of specific circuits affects the result. If necessary, a general introduction to quantum computing can be found in Appendix A. We also recommend the works galvao2024possibilidades; jesus2021computaccao; billig2018quantum; oliveira2021algoritmos; santos2017computador; rabelo2018abordagem; alves2020simulating; jose2013introduccao; perry2019quantum; oliveira2020fisica. Additionally, we suggest materials that may help the reader with specific steps, such as state preparation aulicino2022state and quantum phase estimation alves2022algoritmos; pezze2014quantum.
@@ II.1 Linear Systems of Equations
Linear Systems of Equations (SLE) are fundamental for solving problems in both science and industry. They are widely used to model physical phenomena, such as current flow in electrical circuits, force equilibrium in mechanical structures, and mesh simulation in numerical methods lanczos1952solution; oktacc2018conceptions; crout1941short; kirvan2015flipping. Furthermore, SLEs underpin algorithms in machine learning, optimization, and image processing, and are often solved at large scale on supercomputers cleveland1987progress; saad1989krylov. Thus, developing efficient methods for solving these systems, such as the HHL algorithm in quantum computing, represents an important step toward accelerating scientific and industrial applications.
@@
Here, $A$ is an $N\times N$ matrix and $\vec{x}$ and $\vec{b}$ are $N$-dimensional vectors. In this paper, we adopt the usual notation from Quantum Mechanics, using bras $\left\langle\cdot\right|$ and kets $\left|\cdot\right\rangle$, so that Eq. (1) can be rewritten as $A\left|x\right\rangle=\left|b\right\rangle$ (see Appendix A). Notoriously, the greatest difficulty in solving such systems lies in inverting the matrix $A$ to obtain the solution vector Diblík2016Exponential:
@@
For very large values of $N$, this process requires significant computational resources, since classically the complexity of the problem can scale polynomially zaman. HHL promises to reduce this complexity to logarithmic scale, allowing SLEs to be solved with fewer computational resources harrow2009quantum.
@@ II.2 General Overview of the Algorithm
The HHL algorithm uses in its circuit steps that are widely known in the Quantum Computing literature nielsen2010quantum. Figure 1 shows a general overview of the circuit, explicitly illustrating its routines and applied registers. In summary, in order of implementation, the steps of the algorithm can be synthesized as:
@@ II.4 Quantum Phase Estimation
Quantum Phase Estimation (QPE) was introduced in 1995 by Alexei Kitaev to estimate the eigenvalue phase of an eigenvector associated with a unitary operator kitaev1995quantum. Considering this unitary operator as a matrix $\mathbf{U}$, with eigenvectors $\left|u\right\rangle$, its characteristic equation can be written as
@@ II.4.1 Uniform Superposition
The first step of QPE is similar to the initial stage present in many quantum algorithms: creating a uniformly distributed superposition of states. A famous example of this application is Grover’s algorithm, in which all qubits are initialized this way grover1996fast. In the case of QPE, this superposition is applied only to the registers $q-c:\left|0\right\rangle^{\otimes n}$. This relation can be obtained by applying the Hadamard gate ($H$) to these qubits:
@@ II.4.2 Controlled Unitary Operators
The sequence of products in (16), directly related to $l$ in Eq. (15), is given by the number of qubits that form the qcq-c registers. This number is chosen by the user and defines the precision of the phase calculated by QPE. Using more qubits increases precision, but care must be taken not to incur excessive computational cost for very small changes papadopoulos2024reductive. Moreover, larger circuits are more susceptible to errors due to the current structure of quantum computers. Present-day machines have an intermediate number of physical qubits (between tens and hundreds) and suffer from error and noise rates that prevent the implementation of error correction codes, which characterizes them as NISQ devices (Noisy Intermediate-Scale Quantum) preskill2018quantum. Computational complexity is also a contributing factor in choosing fewer qubits: making the system too complex may waste resources and runtime. A more detailed analysis of the impact of the number of qubits on QPE precision can be found in Ref. Koch2020.
@@
Equation (17) contains a relation known as the Quantum Fourier Transform (QFT), an algorithm used to switch between spaces in quantum systems nielsen2010quantum; coppersmith2002approximate. Its general expression, for two arbitrary states $\left|\alpha\right\rangle$ and $\left|\beta\right\rangle$, can be highlighted as:
@@ II.4.3 Inverse Quantum Fourier Transform
which occurs only when $\theta-y/\tilde{N}=0$, where $\tilde{N}=2^{n}$. This result implies that in any other situation the expression cancels out, since the Fourier transform is based on the sum of sinusoids and, like the Dirac delta function, has a filtering property zaman.
@@ II.5 Ancilla Quantum Encoding
Ancilla qubits are widely used in quantum algorithms, as they assist in specific implementation steps such as indirect measurements garcia2021learning; brida2012ancilla, state control brown2011ancilla; zhang2020pseudogap, or error correction reinhold2020error; criger2012quantum. In the context of HHL, this qubit acts to maximize the probability of obtaining the desired result, given the eigenvalues encoded in the clock qubits during QPE harrow2009quantum. Without loss of generality, one can write a controlled rotation on the state as
@@
Previous works define specific functions to obtain appropriate angles when applying rotations with the gate $R_{Y}(\vec{\theta})$ zaman; zhang2022improved, namely
@@ III Numerical Implementation
Using these values to determine the corresponding rotations in the quantum circuit will be discussed in later subsections. However, this is not a mandatory requirement of the algorithm but rather an efficient didactic choice to understand the theoretical foundations discussed in Sec. II. In this way, one can visualize how specific rotations influence the final result. This is because the HHL algorithm does not require prior knowledge of the eigenvalues and eigenvectors of the matrix $A$; such information is accessed indirectly through quantum phase estimation, provided the matrix is sparse and well-conditioned, and an efficient oracle exists to simulate $e^{iAt}$. Thus, in cases where the matrix $A$ is unitary and can be implemented via quantum logic gates, the appropriate choice of rotations can be made efficiently, without the need for explicit diagonalization harrow2009quantum.
In the following subsections, we present code for implementing the HHL routines using Qiskit, IBM’s open-source framework for programming and simulating quantum circuits on real hardware and simulators; the code can be found in HHL_Quantum_Algorithm. The implementation is carried out in Python, requiring basic familiarity with the language. For an introduction to Python, we recommend dos2022uso; borges2014python.
@@ III.2 Quantum Phase Estimation
$\begin{split}\left|\varphi_{2}\right\rangle&=\left|1\right\rangle\frac{1}{2}(\left|00\right\rangle+\left|01\right\rangle+\left|10\right\rangle+\left|11\right\rangle)\left|0\right\rangle_{a},\\ \left|\varphi_{2}\right\rangle&=\frac{1}{2\sqrt{2}}(-\left|u_{1}\right\rangle\left|00\right\rangle-\left|u_{1}\right\rangle\left|01\right\rangle-\left|u_{1}\right\rangle\left|10\right\rangle-\left|u_{1}\right\rangle\left|11\right\rangle\\ &+\left|u_{2}\right\rangle\left|00\right\rangle+\left|u_{2}\right\rangle\left|01\right\rangle+\left|u_{2}\right\rangle\left|10\right\rangle+\left|u_{2}\right\rangle\left|11\right\rangle)\,{\left|0\right\rangle_{a}}.\end{split}$ (44)
@@
$\mathbf{U}=\frac{-i}{2}\begin{pmatrix}-1+i&1+i\\ 1+i&-1+i\\ \end{pmatrix}\,$, (48)
where i-i corresponds to a global phase factor $e^{-i\pi/2}$, which can be ignored as it does not affect physical observables.
Obtaining higher powers of the unitary reduces to $\mathbf{U^{l}}=V\mathbf{U_{d}^{l}}V^{-1}$. For $n=2$ c-qubits, $l=2$ is sufficient to provide acceptable precision for the algorithm (QPE is a phase estimation and using more qqcc qubits refines this estimate; the measured value gets closer to the expected phase, but the measurement probability may decrease Koch2020). Thus, $\mathbf{U^{2}}$ is written as
@@ III.5 Numerical Results
To further evaluate the algorithm’s potential on NISQ-era hardware preskill2018quantum, we also executed it on a real quantum computer, ibm_kiyv, available via the IBM Quantum Experience; its specifications are shown in Tab. 1. For a more reliable metric, we ran the algorithm $10$ times on this backend and computed the mean and standard deviation. Both results are shown in Fig. 5, with the probability distributions produced by the algorithm.
@@ IV Noise Analysis in the Algorithm
In the context of the NISQ era, several factors can affect the efficiency of the algorithm, such as gate errors, relaxation noise ($T_{1}$), dephasing noise ($T_{2}$), and errors inherent to the measurement process preskill2018quantum. Each of these noise sources contributes in different ways to information loss and to the deviation between ideal and experimentally obtained results.
@@
This behavior highlights the importance of precise control over two-qubit gates—which are notoriously noisier than single-qubit operations—as well as the need to minimize circuit depth to reduce decoherence effects. To mitigate the errors observed on real quantum hardware, several strategies can be applied. One of the most direct approaches is readout error mitigation, which calibrates measurement probabilities using known reference states, enabling a more accurate reinterpretation of the observed outcomes hicks2022active; doi:10.1126/sciadv.abi8009; PhysRevA.106.012423.
@@
In more advanced scenarios, implementing quantum error-correcting codes such as the surface code may be considered, although they still require a large number of physical qubits to protect a single logical qubit, which makes them impractical for many present-day systems PhysRevA.83.020302; Horsman_2012; doi:10.1126/sciadv.1500707. Finally, statistical repetition and post-processing techniques are also useful for smoothing noise effects, especially when combined with theoretical modeling maciejewski2020mitigation; zhukov2022quantum. Simulation with noise models, as presented in this work, also allows one to anticipate behaviors and adjust strategies before running on hardware, representing an essential step toward developing more robust quantum algorithms.